OpenStudy (anonymous):

Solve for x: [3x+8]=0

OpenStudy (anonymous):

x=-8/3

OpenStudy (anonymous):

transpose 8 to right hand side 3x = 0-8 3x = -8 divide both sides by 3 3x/3 = -8/3 x = -8/3

OpenStudy (anonymous):

[3x+8]=0 [brackets not important] 3x+8=0 subtract 8 from both sides 3x=-8 [divide by 3 on both sides] x = -8/3

OpenStudy (anonymous):

how do you guys know the brackets dont mean this? http://www.mathnstuff.com/math/spoken/here/1words/g/g6.htm

OpenStudy (anonymous):

I've something those Brackets Can mean Greatest Integer Function !

OpenStudy (anonymous):

yes joe but here brackets were not important according to me

OpenStudy (anonymous):

squeaky, please say what those brackets are, if you dont, these answer could possible be wring, and we wont know.

OpenStudy (anonymous):

[3x+8]=0 is not the same as |3x+8| which means the absolute value

OpenStudy (anonymous):

wrong*

OpenStudy (anonymous):

I'm with Joemath

OpenStudy (anonymous):

brackets will be useful if e.g you have [3x+8]2 = 0 i guess.. :p

OpenStudy (anonymous):

please look at that link guys, [x] could possible mean something else. im not saying you guys are wrong. We just dont know unless squeaky can say for sure that those are irrelevant brackets or not

OpenStudy (anonymous):

yes yes u r right joe Squeaky u must tell that what the brackets mean here